OpenStudy (anonymous):

Which of the following is the slope between the two points (1, 5) and (6, 4)? -5 negative 1 over 5 5 1 over 5

OpenStudy (anonymous):

slope = change in X / change in Y (6 - 1) ------ (4 - 5) = 5 / -1 = - 5

OpenStudy (anonymous):

sorry I mean, slope = change in Y / change in X (4 - 5) ------ (6 - 1) = -1 / 5

OpenStudy (anonymous):

Yup. The slope formula is: \[m = \frac{y_{1} - y_{2}}{x_{1} - x_{2}}\] Make sure when using this formula, you use the x and y-values from each point in the same order. \[m = \frac{(5 - 4)}{(1 - 6)} = -\frac{1}{5}\]

OpenStudy (anonymous):

See how 5 and 1 are from the same coordinate and they were used in the same order?